# AudioMoG: Guiding Audio Generation with Mixture-of-Guidance

### v1

@@ 6 Conclusion
In this work, we introduce AudioMoG, an improved sampling framework for audio generation. We start from an analysis of CFG and AG methods, demonstrating their respective principle on outperforming unguided diffusion generation. Then, we propose the mixture-of-guidance framework, introducing its mechanism which exploits the cumulative benefits of diverse guidance methods. Comprehensive experiments demonstrate the superiority of AudioMoG over the popularly adopted CFG on T2A generation under the same inference budgets, as well as achieving improvement over both CFG and AG across V2A, T2M, and image generation tasks without increasing inference time.
Appendix A Proof of AudioMoG
@@
$=$ $\{\bm{\epsilon}|\exists\hat{w}_{1},\hat{w}_{2}\in\mathbb{R},\hat{w}_{3}\notin\{0,1\},\text{s.t.}\bm{\epsilon}=\hat{\bm{\epsilon}}_{\text{HG}}(\bm{z}_{t},t,\bm{c})\}$
Proof. Let the two sets in Equation 13 be denoted as $S_{1}$ and $S_{2}$. Substituting the first two equations of Equation 11 into the third gives:
@@
$+w_{3}(1-w_{1})\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})+(1-w_{2})(1-w_{3})\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t)$
Likewise, Equation 12 gives
@@
$+\hat{w}_{2}(1-\hat{w}_{3})\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})+(1-\hat{w}_{2})(1-\hat{w}_{3})\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t)$
For most cases $\bm{\epsilon}_{\theta}(\bm{z}_{t},t)$, $\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})$, $\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t)$ and $\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})$ are linearly independent, thus Equation 14 and 15 implies
@@
For Equations 16, adding the first and third equations yields
$w_{3}=\hat{w}_{1}\hat{w}_{3}+\hat{w}_{2}(1-\hat{w}_{3})$ (17)
@@
Therefore, $S_{2}\subseteq S_{1}$. Notice that Equations 16 is symmetric under the exchange of $w_{1}$ and $\hat{w}_{1}$, $w_{2}$ and $\hat{w}_{2}$, $w_{3}$ and $\hat{w}_{3}$, so similarly
@@
Appendix B Impact of guidance scale
@@
The PG setting. We study the guidance scales $w_{1}$ in the range $4.3\sim 5.0$ and $w_{2}$ in $0.0\sim 0.3$. We keep the NFE fixed at 300 following the main paper and evaluated FAD, IS and CLAP. The results are shown in Figures 5(a)-5(c). Noticing that PG degrades to CFG-only when $w_{2}=0$, we can see that PG indeed reaches a higher performance ceiling and achieves the best results around our choice, $w_{1}=4.6$ and $w_{2}=0.2$. Excessively increasing or decreasing $w_{1}$ and $w_{2}$ will result in a loss of quality, showing alike convex pattern in the following HG setting.
The HG setting. We adopt the setting in the main paper, with the baseline configuration of $w_{1}=4.0$, $w_{2}=3.3$, and $w_{3}=1.2$, while keeping the NFE fixed at 400. We evaluated both FAD and IS as primary metrics, and the results are summarized in Figures 6(a)-6(c). For $w_{1}$ (Figure 6(a)), we can see as $w_{1}$ increases from 3.0 to 4.0, FAD decreases substantially, reaching a minimum around $w_{1}=4.0$, indicating improved overall sample fidelity. However, further increasing $w_{1}$ beyond this point slightly worsens FAD. Meanwhile, IS increases steadily and saturates at $w_{1}\geq 4.0$, reflecting stronger condition adherence and sample specificity. For $w_{2}$ and $w_{3}$, we observe similar convex patterns (Figure 6(b), 6(c)). Both FAD and IS achieve their best values around $w_{2}=3.3$ and $w_{3}=1.2$. Therefore, all three scales exhibit non-monotonic behaviors, and each has an optimal value range where both fidelity and diversity metrics are simultaneously optimized. These results validate the necessity of tuning each guidance component, as also noted in Karras et al. (2024a), and highlight that balanced guidance from multiple perspectives is crucial for high-quality audio generation.
@@ Appendix C Text-to-music generation results
Table 6: Objective metrics for text-to-music generation on MusicCaps dataset.
@@
Riffusion (Forsgren & Martiros, 2022) 13.40 / /
Mubert (Mubert-Inc., 2022) 9.60 / /
MusicLM (Agostinelli et al., 2023) 4.00 / /
Mousai (Schneider et al., 2023) 7.50 / /
MeLoDy (Lam et al., 2023) 5.41 / /
Stable Audio Open (Evans et al., 2025) 3.51 2.93 36.42
MusicGen w/o melody (Copet et al., 2023) 3.40 / /
AudioLDM 2-Large (Liu et al., 2024a) 2.93 2.59 16.34
AudioLDM 2-Full (Liu et al., 2024a) 3.13 / /
TANGO-AF (Kong et al., 2024) 2.21 2.79 22.69
Jen-1 (Li et al., 2024b) 2.00 / /
@@ C.1 Baseline methods
Riffusion (Forsgren & Martiros, 2022) is a unique model that generates music by converting spectrogram images into audio. It fine-tunes the Stable Diffusion model on spectrograms, allowing it to produce short music loops based on text prompts.
Mubert (Mubert-Inc., 2022) is an AI-driven music generation platform that creates royalty-free music tailored for various content needs. It offers tools for content creators, artists, and developers to generate and integrate AI-generated music into their projects.
MusicLM (Agostinelli et al., 2023) is a model introduced by Google that generates high-fidelity music from text descriptions. It utilizes a sequence-to-sequence modeling approach to capture long-term structure in music generation.
Mousai (Schneider et al., 2023) is a two-stage latent-diffusion system for text-to-music generation. Stage 1 compresses 48 kHz stereo audio with a Diffusion-Magnitude Autoencoder (DMAE), and Stage 2 is a text-conditioned latent diffusion (TCLD) model that can produce multi-minute, prompt-aligned musical pieces.
MeLoDy (Lam et al., 2023) is an LM-guided diffusion framework that inherits the highest-level LM from MusicLM for semantic modeling, and applies a novel dual-path diffusion (DPD) model and an audio VAE-GAN to efficiently decode the conditioning semantic tokens into waveform, resulting in cutting sampling cost by more than 95 % while maintaining state-of-the-art text–music alignment and audio quality.
MusicGen (Copet et al., 2023) is an open-source model developed by Meta that generates music from text prompts. It employs a transformer-based architecture trained on a large dataset of music to produce diverse and high-quality audio samples.
Jen-1 (Li et al., 2024b) is a universal high-fidelity model for text-to-music generation. It incorporates both autoregressive and non-autoregressive training, enabling tasks like text-guided music generation, inpainting, and continuation.
C.2 Experiment results
We conduct a comprehensive analysis of generated music quality across the above systems. For Riffusion, Mubert and MusicLM, we report the metrics from MusicLM. For Stable Audio Open, AudioLDM2-Large and Tango-AF, we cite the results from ETTA. For Mousai, as it is not evaluated in ETTA, we report it from AudioLDM 2. For other baselines, we report the metrics as presented in their original papers. Similar to T2A, we further evaluate our baseline model using only CFG, which also achieves the best result when $w=7$, and fix NFE to 400. All evaluations are conducted on MusicCaps dataset. The results are detailed in Table 6. It can be shown that HG not only surpasses CFG-only but also achieves SOTA in all metrics.
@@ Appendix D Image generation results
This section presents the experimental results for the conditional ImageNet $512\times 512$ generation task, using the EDM2-S checkpoints from (Karras et al., 2024b). Image generation was performed over 16 deterministic steps with a second-order Heun sampler, maintaining the same inference speed with a single guidance method (Karras et al., 2024a). The optimal guidance strengths were determined through a grid search on a reduced sample size ($N=8192$). Subsequently, the model’s performance was evaluated on a larger set of samples ($N=50000$) to obtain robust estimates of the Fréchet Inception Distance (FID) and the Fréchet DINOv2 Distance (FD_DINOv2).
Table 7: Conditional image generation results on ImageNet-512.
@@
This section outlines the setup of the 2D toy dataset experiment used in the analysis presented in Section 3.1. Unless otherwise specified, the experiment details strictly follow the setup in (Karras et al., 2024a).
@@
Models. We employ simple multi-layer perceptrons (MLPs) as denoiser models, consistent with the setup in (Karras et al., 2024a).
Training. To ensure comparability, we use the pre-trained models provided by (Karras et al., 2024a). The URLs of model checkpoints can be found at: https://github.com/NVlabs/edm2/blob/main/toy_example.py.
Sampling. For the visualizations in Figure 2, we use the following guidance weights: $w=3$ for both CFG and AG; $w_{1}=2,w_{2}=2$ for PG; and $w_{1}=w_{2}=1.5$, $w_{3}=2$ for HG. The models $\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})$ and $\bm{\epsilon}_{\theta}(\bm{z}_{t},t)$ use checkpoints trained with hidden dimension $d=64$ and training iteration $M=4096$, while the bad models $\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})$ and $\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t)$ use checkpoints with $d=32$ and $M=512$.
@@ Appendix F Text-to-audio experiment details
F.1 Datasets
We present the datasets used to train our baseline model in Table 8. AudioCaps (Kim et al., 2019) is a benchmark dataset for audio captioning that contains 50,000 audio clips from AudioSet paired with human-written textual descriptions. We only used its training set. AudioSet (Gemmeke et al., 2017) is a large-scale weakly labeled dataset released by Google, comprising over 2 million 10-second audio clips across more than 600 sound event categories. The BBC Sound Effects Library [link] provides over 30,000 professionally recorded sound effects covering a wide variety of acoustic scenes and events. Clotho v2 (Drossos et al., 2020) is designed for audio captioning tasks and contains approximately 5,000 audio clips, each annotated with five crowdsourced textual descriptions. VGGSound (Chen et al., 2020) is a large-scale dataset consisting of over 200,000 10-second video clips from YouTube, covering 310 diverse sound classes. FreeSound [link] is a collaborative platform that hosts a wide range of user-contributed audio samples, frequently used for environmental sound classification and retrieval. FSD50K (Fonseca et al., 2021) is a large-scale dataset derived from FreeSound, containing over 50,000 audio clips annotated with strong and weak labels for sound event detection. FMA (Defferrard et al., 2016), the Free Music Archive dataset, includes full-length high-quality music tracks and is widely used in music information retrieval research. The Million Song Dataset (MSD) (Bertin-Mahieux et al., 2011) offers metadata and pre-computed audio features for one million popular music tracks to support large-scale music recommendation and analysis. MagnaTagATune (MTT) (Law et al., 2009) is a music tagging dataset that contains 25,000 audio clips annotated with multiple descriptive tags for genre, instrument, and mood classification tasks.
Table 8: Statistics for the datasets used in the paper.
@@ F.1 Datasets
Clotho v2 152 (Drossos et al., 2020)
VGGSound 550 (Chen et al., 2020)
@@
FSD50k 108 (Fonseca et al., 2021)
FMA 900 (Defferrard et al., 2016)
MSD 7333 (Bertin-Mahieux et al., 2011)
MTT 200 (Law et al., 2009)
F.2 Baseline methods
@@
AudioLDM is a latent diffusion model for text-to-audio (T2A) generation presented by (Liu et al., 2023). It performs the diffusion process in the latent space of a pretrained audio VAE, while conditioning on text embeddings produced by the CLAP text branch. This design enables a T2A training process without paired data and produces audio that is semantically consistent with the input description.
@@
AudioLDM2 is an improved version of the AudioLDM model, introduced by (Liu et al., 2024a). It incorporates several improvements, including leveraging the Language of Audio (LOA) encoder and finetuning a GPT-2 model to translate any modality to LOA. These improvements result in higher-quality audio generation that better aligns with the input text.
@@
AudioGen is a text-to-audio generation model introduced by (Kreuk et al., 2022). It employs an autoregressive transformer architecture to generate audio samples conditioned on textual descriptions. The model is trained on a large-scale dataset of audio-text pairs, enabling it to produce high-quality audio outputs that align with the given textual input.
@@
Make-An-Audio is a diffusion-based text-to-audio generation model proposed by (Huang et al., 2023b). It introduces a pseudo prompt enhancement with the distill-then-reprogram approach, including a large number of concept compositions by opening up the usage of language-free audios to alleviate data scarcity. Therefore, it enables high-fidelity, prompt-aligned outputs.
@@
TANGO-AF&AC-FT-AC (Kong et al., 2024) pre-trains the TANGO architecture on the synthetic-caption AF-AudioSet plus AudioCaps, followed by fine-tuning on AudioCaps alone. Leveraging high-quality synthetic captions significantly improves text-to-audio alignment and overall audio realism.
@@
Stable Audio Open (Evans et al., 2025) is an open-source text-to-audio generation model developed by Stability AI. It leverages a diffusion-based architecture trained on a diverse dataset of audio-text pairs. The model is designed to generate high-fidelity audio samples conditioned on textual input, supporting various applications such as music generation, sound effect synthesis, and more.
F.3 Model configurations

### v2

@@ VI Conclusion
In this work, we introduce AudioMoG, a mixture-of-guidance framework to improve audio generation quality in the sampling process. In particular, we present two mixture strategies, PG and HG, which directly combine complementary guidance methods, or hierarchically integrate them for cumulative improvements. Comprehensive experiments demonstrate the superiority of AudioMoG over the popularly adopted CFG on T2A generation under the same inference budgets, as well as achieving improvement over both CFG and AG across V2A, T2M, and image generation tasks without increasing inference time.
Appendix A Proof of HG
@@
$=$ $\{\bm{\epsilon}|\exists\hat{w}_{1},\hat{w}_{2}\in\mathbb{R},\hat{w}_{3}\notin\{0,1\},\text{s.t.}\bm{\epsilon}=\hat{\bm{\epsilon}}_{\text{HG}}(\bm{z}_{t},t,\bm{c})\}.$
Proof. Let the two sets in (12) be denoted as $S_{1}$ and $S_{2}$. Substituting the first two equations of (10) into the third gives:
@@
$+w_{3}(1-w_{1})\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})+(1-w_{2})(1-w_{3})\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t).$
Likewise, Equation (11) gives
@@
$+\hat{w}_{2}(1-\hat{w}_{3})\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})+(1-\hat{w}_{2})(1-\hat{w}_{3})\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t).$
For most cases $\bm{\epsilon}_{\theta}(\bm{z}_{t},t)$, $\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})$, $\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t)$ and $\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})$ are linearly independent, thus Equations (13) and (14) imply
@@
For Equations (15), adding the first and third equations yields
$w_{3}=\hat{w}_{1}\hat{w}_{3}+\hat{w}_{2}(1-\hat{w}_{3}).$ (16)
@@
Therefore, $S_{2}\subseteq S_{1}$. Notice that Equations (15) is symmetric under the exchange of $w_{1}$ and $\hat{w}_{1}$, $w_{2}$ and $\hat{w}_{2}$, $w_{3}$ and $\hat{w}_{3}$, so similarly
@@
Appendix B Impact of guidance scales
@@
The PG setting. We study the guidance scales $w_{1}$ in the range $4.3\sim 5.0$ and $w_{2}$ in $0.0\sim 0.3$. We keep the NFE fixed at 300 following the main paper and evaluated FAD, IS and CLAP. The results are shown in Fig. 4(a)-4(c). Noticing that PG degrades to CFG-only when $w_{2}=0$, we can see that PG indeed reaches a higher performance ceiling and achieves the best results around our choice, $w_{1}=4.6$ and $w_{2}=0.2$. Excessively increasing or decreasing $w_{1}$ and $w_{2}$ will result in a loss of quality, showing alike convex pattern in the following HG setting.
The HG setting. We adopt the setting in the main paper, with the baseline configuration of $w_{1}=4.0$, $w_{2}=3.3$, and $w_{3}=1.2$, while keeping the NFE fixed at 400. We evaluated both FAD and IS as primary metrics, and the results are summarized in Fig. 5(a)-5(c). For $w_{1}$ (Fig. 5(a)), we can see as $w_{1}$ increases from 3.0 to 4.0, FAD decreases substantially, reaching a minimum around $w_{1}=4.0$, indicating improved overall sample fidelity. However, further increasing $w_{1}$ beyond this point slightly worsens FAD. Meanwhile, IS increases steadily and saturates at $w_{1}\geq 4.0$, reflecting stronger condition adherence and sample specificity. For $w_{2}$ and $w_{3}$, we observe similar convex patterns (Fig. 5(b), 5(c)). Both FAD and IS achieve their best values around $w_{2}=3.3$ and $w_{3}=1.2$. Therefore, all three scales exhibit non-monotonic behaviors, and each has an optimal value range where both fidelity and diversity metrics are simultaneously optimized. These results validate the necessity of tuning each guidance component, as also noted in [31], and highlight that balanced guidance from multiple perspectives is crucial for high-quality audio generation.
@@ Appendix C Text-to-music generation results
TABLE V: Objective metrics for text-to-music generation on MusicCaps dataset.
@@
Riffusion [16] 13.40 / /
Mubert [49] 9.60 / /
MusicLM [1] 4.00 / /
Mousai [52] 7.50 / /
MeLoDy [36] 5.41 / /
Stable Audio Open [14] 3.51 2.93 36.42
MusicGen w/o melody [7] 3.40 / /
AudioLDM 2-Large [43] 2.93 2.59 16.34
AudioLDM 2-Full [43] 3.13 / /
TANGO-AF [34] 2.21 2.79 22.69
Jen-1 [40] 2.00 / /
@@ C-A Baseline methods
Riffusion [16] is a unique model that generates music by converting spectrogram images into audio. It fine-tunes the Stable Diffusion model on spectrograms, allowing it to produce short music loops based on text prompts.
Mubert [49] is an AI-driven music generation platform that creates royalty-free music tailored for various content needs. It offers tools for content creators, artists, and developers to generate and integrate AI-generated music into their projects.
MusicLM [1] is a model introduced by Google that generates high-fidelity music from text descriptions. It utilizes a sequence-to-sequence modeling approach to capture long-term structure in music generation.
Mousai [52] is a two-stage latent-diffusion system for text-to-music generation. Stage 1 compresses 48 kHz stereo audio with a Diffusion-Magnitude Autoencoder (DMAE), and Stage 2 is a text-conditioned latent diffusion (TCLD) model that can produce multi-minute, prompt-aligned musical pieces.
MeLoDy [36] is an LM-guided diffusion framework that inherits the highest-level LM from MusicLM for semantic modeling, and applies a novel dual-path diffusion (DPD) model and an audio VAE-GAN to efficiently decode the conditioning semantic tokens into waveform, resulting in cutting sampling cost by more than 95 % while maintaining state-of-the-art text–music alignment and audio quality.
MusicGen [7] is an open-source model developed by Meta that generates music from text prompts. It employs a transformer-based architecture trained on a large dataset of music to produce diverse and high-quality audio samples.
Jen-1 [40] is a universal high-fidelity model for text-to-music generation. It incorporates both autoregressive and non-autoregressive training, enabling tasks like text-guided music generation, inpainting, and continuation.
C-B Experiment results
We conduct a comprehensive analysis of generated music quality across the above systems. For Riffusion, Mubert and MusicLM, we report the metrics from MusicLM. For Stable Audio Open, AudioLDM2-Large and Tango-AF, we cite the results from ETTA. For Mousai, as it is not evaluated in ETTA, we report it from AudioLDM 2. For other baselines, we report the metrics as presented in their original papers. Similar to T2A, we further evaluate our baseline model using only CFG, which also achieves the best result when $w=7$, and fix NFE to 400. All evaluations are conducted on MusicCaps dataset. The results are detailed in Table V. It can be shown that HG not only surpasses CFG-only but also achieves SOTA in all metrics.
@@ Appendix D Image generation results
[…]
TABLE VI: Conditional image generation results on ImageNet-512.
@@
This section outlines the setup of the 2D toy dataset experiment used in the analysis presented in Section III. Unless otherwise specified, the experiment details strictly follow the setup in [31].
@@
Models. We employ simple multi-layer perceptrons (MLPs) as denoiser models, consistent with the setup in [31].
Training. To ensure comparability, we use the pre-trained models provided by [31]. The URLs of model checkpoints can be found at: https://github.com/NVlabs/edm2/blob/main/toy_example.py.
Sampling. For the visualizations in Fig. 2, we use the following guidance weights: $w=3$ for both CFG and AG; $w_{1}=2,w_{2}=2$ for PG; and $w_{1}=w_{2}=1.5$, $w_{3}=2$ for HG. The models $\bm{\epsilon}_{\theta}(\bm{z}_{t},t,\bm{c})$ and $\bm{\epsilon}_{\theta}(\bm{z}_{t},t)$ use checkpoints trained with hidden dimension $d=64$ and training iteration $M=4096$, while the bad models $\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t,\bm{c})$ and $\bm{\epsilon}_{\theta_{\text{bad}}}(\bm{z}_{t},t)$ use checkpoints with $d=32$ and $M=512$.
@@ Appendix F Text-to-audio experiment details
F-A Datasets
We present the datasets used to train our baseline model in Table VII. AudioCaps [33] is a benchmark dataset for audio captioning that contains 50,000 audio clips from AudioSet paired with human-written textual descriptions. We only used its training set. AudioSet [17] is a large-scale weakly labeled dataset released by Google, comprising over 2 million 10-second audio clips across more than 600 sound event categories. The BBC Sound Effects Library [link] provides over 30,000 professionally recorded sound effects covering a wide variety of acoustic scenes and events. Clotho v2 [11] is designed for audio captioning tasks and contains approximately 5,000 audio clips, each annotated with five crowdsourced textual descriptions. VGGSound [4] is a large-scale dataset consisting of over 200,000 10-second video clips from YouTube, covering 310 diverse sound classes. FreeSound [link] is a collaborative platform that hosts a wide range of user-contributed audio samples, frequently used for environmental sound classification and retrieval. FSD50K [15] is a large-scale dataset derived from FreeSound, containing over 50,000 audio clips annotated with strong and weak labels for sound event detection. FMA [10], the Free Music Archive dataset, includes full-length high-quality music tracks and is widely used in music information retrieval research. The Million Song Dataset (MSD) [2] offers metadata and pre-computed audio features for one million popular music tracks to support large-scale music recommendation and analysis. MagnaTagATune (MTT) [37] is a music tagging dataset that contains 25,000 audio clips annotated with multiple descriptive tags for genre, instrument, and mood classification tasks.
TABLE VII: Statistics for the datasets used in the paper.
@@ F-A Datasets
Clotho v2 152 [11]
VGGSound 550 [4]
@@
FSD50k 108 [15]
FMA 900 [10]
MSD 7333 [2]
MTT 200 [37]
F-B Baseline methods
@@
AudioLDM is a latent diffusion model for text-to-audio (T2A) generation presented by [42]. It performs the diffusion process in the latent space of a pretrained audio VAE, while conditioning on text embeddings produced by the CLAP text branch. This design enables a T2A training process without paired data and produces audio that is semantically consistent with the input description.
@@
AudioLDM2 is an improved version of the AudioLDM model, introduced by [43]. It incorporates several improvements, including leveraging the Language of Audio (LOA) encoder and finetuning a GPT-2 model to translate any modality to LOA. These improvements result in higher-quality audio generation that better aligns with the input text.
@@
AudioGen is a text-to-audio generation model introduced by [35]. It employs an autoregressive transformer architecture to generate audio samples conditioned on textual descriptions. The model is trained on a large-scale dataset of audio-text pairs, enabling it to produce high-quality audio outputs that align with the given textual input.
@@
Make-An-Audio is a diffusion-based text-to-audio generation model proposed by [23]. It introduces a pseudo prompt enhancement with the distill-then-reprogram approach, including a large number of concept compositions by opening up the usage of language-free audios to alleviate data scarcity. Therefore, it enables high-fidelity, prompt-aligned outputs.
@@
TANGO-AF&AC-FT-AC [34] pre-trains the TANGO architecture on the synthetic-caption AF-AudioSet plus AudioCaps, followed by fine-tuning on AudioCaps alone. Leveraging high-quality synthetic captions significantly improves text-to-audio alignment and overall audio realism.
@@
Stable Audio Open [14] is an open-source text-to-audio generation model developed by Stability AI. It leverages a diffusion-based architecture trained on a diverse dataset of audio-text pairs. The model is designed to generate high-fidelity audio samples conditioned on textual input, supporting various applications such as music generation, sound effect synthesis, and more.
F-C Model configurations